# ACT: Agentic Classification Tree

### v2

@@ 4 Experiments
Models. We evaluate our method on four language models: Gemma-4B Team et al. (2025), GPT-Nano-4.1 Achiam et al. (2023), GPT-Mini-4.1 Achiam et al. (2023), and Qwen3-4B Yang et al. (2025). These models were selected to represent diverse architectural approaches while prioritizing computationally efficient smaller models. To examine the effect of model size, we include two variants of the GPT-4.1 family with different parameter scales.
Baselines. We evaluate ACT against four classification baselines:
  • Chain-of-Thought (CoT, zero-shot) (Wei et al., 2022): standard zero-shot prompting with step-by-step reasoning, without access to labeled examples or demonstrations.
  • DSPy (8-shot in-context learning) (Khattab et al., 2023): a demonstration-based approach that conditions the model on eight curated input–label examples to elicit task-relevant behavior.
  • TextGrad (Yuksekgonul et al., 2024): a prompt optimization method that refines a single task-specific prompt using textual feedback. The prompt includes the task description and class labels, and is optimized for accuracy. We follow the original setup, with no demonstrations provided.
  • TF-IDF + CART: a CART trained on TF-IDF representations of the raw texts (Salton et al., 1975; Breiman et al., 1984) that does not rely on a language model, offering structural transparency but limited semantic interpretability, as it lacks natural-language reasoning.
@@
The three LLM-based baseline methods above (CoT, DSPy, TextGrad) are initialized with task-specific prompts that explicitly state the classification objective and define the target classes (e.g., for DIAGNO, whether a case indicates Tuberculosis). In contrast, ACT begins from a generic, task-agnostic query (“Based on the provided example, does it belong to the positive class?”), without access to the task description, class names, or any domain-specific information. Consequently, ACT must uncover the task through successive node’s question-refinement iterations, making the comparison conservative with respect to considered baselines. We vary two hyperparameters of the proposed ACT algorithm: the maximum tree depth $d$ and the number of prompt refinement steps per node $k$, exploring different configurations in our experiments. To control input length and computational cost, $f_{\text{purity}}$ is restricted to at most $m$ randomly selected well-classified and $m$ randomly selected misclassified instances per group (with $m=50$ in our experiments). To ensure a fair comparison, all methods are evaluated using identical underlying language models.
@@ 4.1 Experimental Results and Discussion
As shown in Table 2, ACT with appropriately selected hyperparameters consistently matches or surpasses the CoT, TextGrad, and DSPy baselines across datasets and LLMs. Concretely, averaged over all datasets and models, the best-performing ACT configuration ($d=4$, $k=20$ ) improves test accuracy by $5.2$ percentage points over DSPy (8-shot) and by $4.7$ points over TextGrad, while providing transparent and interpretable question-based decision paths.
Additionally, we compare our method against CART with TF-IDF features at depth 3, which exhaustively searches for optimal splits based on the TF-IDF matrix. This baseline achieves strong performance (78.8% on DIAGNO, 89.3% on SPAM, 91.3% on JAILBREAK). ACT consistently outperforms TF-IDF+CART on SPAM and JAILBREAK for both GPT-4.1 variants, while only Mini exceeds TF-IDF+CART on DIAGNO. However, the TF-IDF+CART approach remains fundamentally limited by its lack of semantic interpretability. The resulting decision trees (see, e.g., Fig. 5 and 6 in Appendix) are typically hard to understand for users, especially when compared to the natural-language questions learned with ACT (cf. Figure 1 for DIAGNO, Figures 3 and 4 for SPAM and JAILBREAK). This stems in part from TF-IDF+CART selecting single words or n-grams as decision criteria, without any regard to context or semantic meaning. As a result, these models struggle with negation, rephrasing, and adversarially crafted text—features that are particularly prevalent in SPAM and JAILBREAK. These limitations explain ACT’s consistent advantage in these domains. Furthermore, Appendix B.2 compares the most relevant symptoms for Tuberculosis identified by ACT on the DIAGNO dataset with those from established medical sources, highlighting their strong alignment and the efficacy of ACT in high-stakes scenarios.
4.2 Ablation studies
The proposed ACT algorithm has two key hyperparameters: the depth of the tree $d$ and the number of optimization steps per node during training $k$. Additional hyperparameters with less influence include the number $m$ of examples $X_{correct}$ and $X_{error}$ provided to the LLM when performing semantic analysis ($f_{\text{purity}}$) and the maximum number of logical operators $L$ permitted in each generated question (cf. Appendix A.2 for more discussion). As illustrated in Figure 2, the depth $d$ and optimization steps $k$ significantly influence ACT performance. Across datasets and LLMs, test accuracy peaks at depths 4–5, with deeper trees exhibiting decreased test accuracy despite increased training accuracy at depth 6 —an indication of overfitting. Conversely, increasing optimization steps consistently improves performance, though gains plateau after 10–20 steps depending on the configuration, indicating that only a modest number of steps $k$ is sufficient to identify the best splitting question.
Figure 2: Ablation study of ACT hyperparameters on the DIAGNO dataset using GPT-4.1 Nano. (a) Effect of depth parameter $d$ on model accuracy for different numbers of optimization steps per node during training $k$. (b) Effect of the number of optimization steps per node during training $k$ on model accuracy for different depth values $d$. Solid lines represent training accuracy; dashed lines represent test accuracy.
@@ 5 Conclusion
We introduced the Agentic Classification Tree (ACT), a novel framework that combines the interpretability of traditional decision trees with the semantic reasoning capabilities of large language models (LLMs). Unlike conventional decision trees that rely on rigid feature-based splits, ACT dynamically optimizes natural-language prompts at each node, using LLM responses to perform semantically meaningful binary splits. Our experiments on text-based binary classification tasks demonstrate that ACT achieves competitive accuracy while providing interpretable, language-based decision logic. This approach highlights the potential of combining classical machine learning structures with modern language model reasoning, opening new avenues for interpretable and effective decision-making in complex, text-rich domains. Future work includes extending ACT to multi-class classification and regression task, improving computational efficiency through targeted prompt optimization strategies, and exploring applications beyond text classification to other structured and unstructured data modalities.
Appendix A Appendix
A.1 Illustrative Example: Prompt Refinement via Semantic Feedback
@@
We describe the iterative prompt refinement procedure used in ACT, using an illustrative case from the DIAGNO3 medical diagnosis task. At each decision node, the objective is to construct a binary natural-language question that separates the data into semantically meaningful groups, thereby reducing weighted Gini impurity.
At each iteration $k$, the following steps are performed:
@@
    The data is partitioned into two subsets: those routed to the yes branch (denoted sdL) and those routed to the no branch (denoted sdG).
@@
This iterative refinement continues until a stopping criterion is met (e.g., convergence or maximum number of updates). In practice, we observe that prompts evolve from generic task-agnostic formulations to more specific, domain-relevant queries. For instance, an initial prompt such as “Does this example belong to the positive class?” may be refined into a targeted question like “Does the example mention coughing up blood or high fever?”, better aligned with the semantic distinctions in the data.
@@
A.2 Prompt Constraints during Prompt Optimization
During prompt optimization in ACT, additional constraints are applied to keep questions clear and interpretable while encouraging exploration. The goal of these constraints is twofold: (i) ensure that generated questions remain simple, precise, and auditable, and (ii) incentivize diversity so that ACT explores many different semantic formulations during refinement. The maximum number of logical operators $L$ allowed per question is a controllable hyperparameter in ACT.
@@ A.2 Prompt Constraints during Prompt Optimization
The following constraints must always be satisfied: The question has to be clear and easy. The question must focus on at most two characteristics. The question must include at least one characteristic different from the previous one. The content of the new question must be significantly different from the current one. Use at most $L$ logical operators (and/or). The question has to be answerable with yes or no only. The question must finish with “(yes/no)?”. Do not use vague words like could, might, or possibly. Do not use blanks or placeholder tags like “___” or <...>.
A.3 Review of Traditional Decision Tree Methods: CART and C4.5
Classification and Regression Trees (CART), proposed by Breiman et al. (1984), are widely used decision tree algorithms for supervised classification tasks. In this work, we focus exclusively on the classification setting.
@@
C4.5 (Salzberg, 1994) follows the same recursive tree-building process as CART but differs mainly in two respects: (i) it chooses splits by maximizing the information gain ratio rather than minimizing Gini impurity, and (ii) it allows multiway splits on categorical attributes, whereas CART restricts all splits to be binary.
@@
Appendix B Additional results and datasets
B.1 Details on datasets
We consider three publicly available binary text classification datasets spanning different domains:
  • DIAGNO (Ninaa510, 2024): a medical diagnosis dataset (tuberculosis vs. allergic sinusitis) , a derived dataset based on DDXPlus Fansi Tchango et al. (2022). We constructed a balanced subset with 600 training, 100 validation, and 600 test samples, ensuring a 50/50 split (300 tuberculosis and 300 allergic sinusitis cases).
  • SPAM: an email spam detection dataset derived from Deysi (2023). We constructed a balanced subset with 600 training, 100 validation, and 600 test samples, following the same procedure as for DIAGNO.
  • JAILBREAK (Shen et al., 2024): a jailbreak prompt classification dataset with 923 training, 102 validation, and 249 test samples (32 examples have been dropped due to high context tokens).
B.2 Qualitative analysis of the questions generated by ACT

### v3

@@ 4.1 Setup: Datasets, Models & Baselines
Models. We evaluate our method on four language models: Gemma-4B (team2025gemma), GPT-Nano-4.1 (achiam2023gpt), GPT-Mini-4.1 (achiam2023gpt), and Qwen3-4B (yang2025qwen3). These models were selected to represent diverse architectural approaches while prioritizing computationally efficient smaller models. To assess the effect of model scale, we report additional scaling results with full GPT-4.1 (achiam2023gpt).
Baselines. We evaluate ACT against six baselines: Chain-of-Thought (CoT, zero-shot) (wei2022chain). DSPy (BFSR, 8 demos) (khattab2023dspy) following the setup of (yuksekgonul2024textgrad): Bootstrapped Few-Shot Random Search (BFSR) generates reasoning traces and answers for all training samples, filters for correct answers, constructs candidate prompts with up to 8 correct demonstrations (input, reasoning trace and output), evaluates the candidate prompts on the whole training set and selects the best-performing one. TextGrad (yuksekgonul2024textgrad) refines a task-specific prompt through textual feedback, optimizing for accuracy given the task description and class labels. We use a batch size of 50 samples and run 12 epochs with 3 optimization steps each, yielding 36 refinement iterations where each candidate is evaluated on a validation set and the best is retained. TF-IDF + CART trains a CART classifier (depths 3–5) on TF-IDF text representations (salton1975vector; DBLP:books/wa/BreimanFOS84), exhaustively searching for optimal splits without relying on a language model. This offers structural transparency but limited semantic interpretability. The best-performing depth per dataset is reported in Table 2, with full results in Appendix E. LLM Concept Bottleneck + CART (Con. tag.) is a concept-bottleneck baseline where an LLM extracts salient domain concepts and encodes each text as a binary concept vector, on which a shallow CART is then trained. RuleFit (TF-IDF rules) (friedman2008predictive) is a rule-ensemble method that learns sparse, human-readable if–then rules over TF-IDF features.
Configurations.
@@
The three prompt-based baseline methods (CoT, DSPy, TextGrad) are initialized with task-specific prompts that explicitly state the classification objective and define the target classes (e.g., for DIAGNO, whether a case indicates Tuberculosis). In contrast, ACT begins from a generic, task-agnostic query (“Based on the provided example, does it belong to the positive class?”), without access to the task description or class names. Consequently, ACT must uncover the task through successive node’s question-refinement iterations, making the comparison conservative with respect to considered baselines. We vary two hyperparameters of the proposed ACT algorithm: the tree depth $d$ and the number of prompt refinement steps per node $k$. To control input length and computational cost, $f_{\text{purity}}$ uses at most $m$ randomly selected well-classified and $m$ randomly selected misclassified instances per group (with $m=50$ in our experiments).
@@ 4.2 Experimental Results and Discussion
As shown in Table 2, ACT with appropriately selected hyperparameters consistently matches or surpasses the four LLM-based baselines (CoT, TextGrad, Con. Tag. and DSPy) across datasets and LLMs. Averaged over all datasets and LLMs, the best-performing ACT configuration improves test accuracy by $4.7$ percentage points over DSPy and by $5.1$ points over TextGrad, while providing interpretable decision paths. When evaluated by F1-score, ACT outperforms all four LLM-based baselines in 15 of the 20 dataset-model configurations.
Compared to TF-IDF + CART, ACT consistently achieves similar or higher accuracy on all tasks except for DIAGNO, where only the GPT-4.1-mini variant surpasses this baseline. However, TF-IDF+CART remains fundamentally limited by its lack of semantic interpretability. The resulting decision trees (cf. Figures 7 and 8) are typically difficult to interpret compared to the natural-language questions learned by ACT (cf. Figures 15, and 6). This stems from TF-IDF+CART selecting individual words or n-grams as decision criteria without regard to context or semantic meaning, causing the model to struggle with negation, irony, rephrasing and adversarially crafted text—features particularly prevalent in SPAM, JAILBREAK, and IMDB. Compared to RuleFit, ACT with GPT-4.1-mini consistently achieves higher accuracy on all tasks, but RuleFit outperforms ACT on average for DIAGNO and JAILBREAK. Like TF-IDF + CART, RuleFit offers limited interpretability because its rules operate on TF-IDF embeddings (e.g., nasal $\leq 0.238$), whereas ACT provides human-readable yes/no questions. Additional experiments (Appendix E) show that ACT remains competitive with strong non-interpretable baselines (XGBoost, BERT, RoBERTa) while providing interpretable and human-readable decision paths.
Table 3: Interpretability Statistics averaged over the 4 LLMs evaluated on DIAGNO. For reference, the generic seed question “Based on … positive class? (yes/no)” has a length of 87 characters.
ACT ($d$, $k$) Avg. Q. Len Avg. Path
(3, 10) 92.00 2.35
(3, 20) 93.15 2.51
(4, 10) 78.05 3.03
(4, 20) 80.53 3.06
Interpretability.
To quantify ACT’s interpretability, we report the average path length at inference and average node-question length in Table 3 (full results in Appendix F). Predictions require only about three nodes even for depth-4 trees and questions remain easily understandable ($\leq$ 155 characters), indicating that soft constraints on question complexity, particularly the limit $L$ on logical clauses, are effective. These quantitative results, together with manual inspection of generated trees (Figure 1; Appendix D.2), confirm that ACT’s divide-and-conquer approach of decomposing complex classification tasks into simpler subquestions yields interpretable decision paths while improving accuracy. Furthermore, Table 4 compares the symptoms for Tuberculosis identified by ACT with established medical sources  (who2024tuberculosis), revealing strong alignment and demonstrating ACT’s efficacy in high-stakes scenarios (cf. Appendix D.1).
Table 4: Symptoms comparison between ACT and medical sources.
In common (5): Coughing up blood, Fever, Weight loss, Swollen lymph nodes, Fatigue
ACT only (2): Shortness of breath, Severe pain
Medical sources only (4): Chest pain, Headache, Increased sweating, Loss of appetite
Figure 2: Ablation study of ACT hyperparameters on DIAGNO using GPT-4.1 Nano. Top: effect of tree depth $d$. Bottom: effect of optimization steps $k$.
Ablation studies.
As shown in Figure 2, tree depth $d$ and optimization steps per node $k$ both significantly influence ACT’s performance. As expected for tree-based methods, increasing depth beyond a certain point (e.g., from 5 to 6 for DIAGNO with GPT-4.1-Nano) leads to overfitting. Increasing $k$ consistently improves performance, though gains plateau after 10–20 steps. Additional results, including ablations on the number $m$ of examples provided to $f_{\text{purity}}$ during semantic analysis, are in Appendix G.
Figure 3: Scaling analysis across the GPT-4.1 family. Test accuracy (%) versus training cost (USD) for TextGrad, DSPy, ACT $(d=3,k=10)$ and ACT $(d=3,k=20)$ on DIAGNO using Azure-hosted GPT-4.1 models.
Scaling analysis.
We evaluate how ACT performance scales with model capacity across the GPT-4.1 family (Nano, Mini and full GPT-4.1). As shown in Figure 3, ACT’s test accuracy consistently improves with stronger models, whereas DSPy and TextGrad do not exhibit consistent scaling behavior, indicating that ACT more effectively leverages the enhanced reasoning capabilities of larger models. This observed scaling trend also suggests a structural advantage of ACT over classical tree-based baselines such as TF-IDF + CART and RuleFit: because ACT can be instantiated with any API-accessible LLM, its performance benefits directly from advances in language model development.
@@ 5 Conclusion
We introduced the Agentic Classification Tree (ACT), a novel framework that combines the interpretability of traditional decision trees with the semantic reasoning capabilities of LLMs. Unlike conventional decision trees that rely on rigid feature-based splits, ACT dynamically optimizes natural-language prompts at each node, using LLM responses to perform semantically meaningful binary splits. Our experiments on text-based binary classification tasks demonstrate that ACT achieves competitive accuracy while providing interpretable, language-based decision logic. This approach highlights the potential of combining classical machine learning structures with modern language model reasoning, opening new avenues for interpretable and effective decision-making in complex, text-rich domains. Future work includes extending ACT to multi-class classification and regression tasks, improving computational efficiency through targeted prompt optimization strategies, and exploring applications beyond text classification to other structured and unstructured data modalities.
Impact Statement
Algorithmic decision making has a broad range of ethical implications, for example, allocative harms or representational harms (shelby2023sociotechnicalharmsalgorithmicsystems). Bias in algorithmic decision making especially often harms minorities (lum-to-predict-and-serve), (pmlr-v81-buolamwini18a), (obermeyer-racial-bias). To reduce these harms and for other reasons, recent governance frameworks and regulations such as the EU AI Act [^2] https://artificialintelligenceact.eu/, the OECD AI Principles [^3] https://oecd.ai/en/ai-principles and the NIST AI Risk Management Framework [^4] https://www.nist.gov/itl/ai-risk-management-framework emphasize that AI-based decisions in high-stakes scenarios must be explainable and subject to human oversight. By providing explicit, interpretable decision paths that are accessible not only to technical experts but also to stakeholders, auditors and regulators, ACT offers an algorithmic decision-making framework that combines competitive accuracy with transparency and contestability.
Nonetheless, two important limitations must be considered. First, while ACT reduces reasoning errors, it remains LLM-based and therefore may still exhibit hallucinations, inconsistencies or unreliability. Second, as noted by (oneil-wmp), transparency and contestability, even if guaranteed by the algorithm, are only effective in practice when embedded within appropriate institutional or regulatory frameworks that provide meaningful avenues for recourse. We therefore urge practitioners deploying ACT in real-world settings to ensure adequate human oversight and mechanisms that allow decisions to be contested in a timely and effective manner.
Appendix A Computational Cost Analysis
Computational cost analysis.
Figure 4 compares the training and inference costs of DSPy, TextGrad and two ACT configurations on the DIAGNO task. We report wall-clock time and total token usage (input plus output). During training, all methods have access to the full training set and at inference time all methods are evaluated on the full test set. All experiments were conducted using Gemma-3-4B-IT on a single H200 GPU with 8 threads per program and results are averaged over five independent runs.
The observed differences reflect the underlying program-construction strategies of each method. TextGrad exhibits relatively high wall-clock time despite lower token usage, as it relies on fewer but substantially longer prompts. DSPy, by contrast, evaluates a larger number of medium-sized prompts derived from few-shot reasoning traces. ACT issues many short prompts to estimate Gini impurity at each node, complemented by a smaller number of longer prompts for semantic feedback and node-level optimization. In its cheapest configuration, ACT is competitive with both baselines in overall computational cost. Increasing tree depth raises training cost but yields higher accuracy while preserving interpretability. At inference time, TextGrad is the most economical, while ACT remains within the same order of magnitude and provides higher accuracy with interpretable decision paths. DSPy shows similar wall-clock time to ACT but incurs higher token usage due to the substantial size of its few-shot prompts.
Figure 4: Training and inference computational cost comparison on DIAGNO. Each plot reports the average wall-clock time and token usage over 5 runs for TextGrad, DSPy (BFSR, 8-shot), and ACT under both the cheapest (d = 3, k = 10) and most expensive (d = 4, k = 20) configurations, using Gemma-3-4B on a single H200 GPU with 8 threads per program. Training cost reflects program construction dynamics: TextGrad processes fewer but very long prompts; DSPy evaluates 13 medium-sized candidate programs derived from few-shot traces; and ACT issues many short node-level prompts for Gini estimation supplemented with a smaller number of larger prompts for semantic feedback and node optimization. Inference cost shows that TextGrad is the most economical, while ACT remains within the same order of magnitude and provides higher accuracy with more interpretable outputs. DSPy exhibits similar wall-clock time to ACT but requires more tokens, largely due to the substantial size of the few-shot prompts.
Appendix B Additional Implementation Details of the ACT Algorithm
B.1 Illustrative Example: Prompt Refinement via Semantic Feedback
@@
We describe the iterative prompt refinement procedure used in ACT, using an illustrative case from the DIAGNO medical diagnosis task. At each decision node, the objective is to construct a binary natural-language question that separates the data into semantically meaningful groups, thereby reducing weighted Gini impurity. At each iteration $k$, the following steps are performed:
@@
    The data is partitioned into two subsets: those routed to the yes branch (denoted $sd_{L}$) and those routed to the no branch (denoted $sd_{G}$).
@@
This iterative refinement continues until a stopping criterion is met (e.g., maximum number of updates). In practice, we observe that prompts evolve from generic task-agnostic formulations to more specific, domain-relevant queries. For instance, an initial prompt such as “Does this example belong to the positive class?” may be refined into a targeted question like “Does the example mention coughing up blood or fever?”, better aligned with the semantic distinctions in the data.
@@
B.2 Prompt Constraints during Prompt Optimization
During prompt optimization in ACT, additional constraints are applied to keep questions clear and interpretable while encouraging exploration. The goal of these constraints is twofold: (i) ensure that generated questions remain simple, precise, and auditable, and (ii) incentivize diversity so that ACT explores many different semantic formulations during refinement. The maximum number of logical operators $L$ allowed per question is a controllable hyperparameter in ACT. In all experiments reported in the paper, we set $L=2$.
@@ B.2 Prompt Constraints during Prompt Optimization
The following constraints must always be satisfied: – The question has to be clear and easy. – The question must include at least one characteristic different from the previous one. – The content of the new question must be significantly different from the current one. – Use at most $L$ logical operators (and/or). – The question has to be answerable with yes or no only. – The question must finish with “(yes/no)?”. – Do not use vague words like could, might, or possibly. – Do not use blanks or placeholder tags like “___” or <...>.
B.3 Details on Datasets
We use five publicly available text classification datasets spanning diverse domains. DIAGNO, SPAM, and JAILBREAK are included because they represent tasks for which structured reasoning is expected to be beneficial, aligning with the types of problems ACT is designed to address. BANKCHURN, a tabular dataset serialized into text, is included because it relies less on pretraining-aligned task semantics and more on learning dataset-specific relationships between attributes and the churn label. IMDB, a widely used benchmark for sentiment classification, is added both to broaden the evaluation and to facilitate comparison with prior work. All datasets are relatively small and mostly balanced, reflecting the low-data prompt-optimization regime in which ACT and comparable methods are typically evaluated (yuksekgonul2024textgrad).
  • DIAGNO (ninaa510diagnosis): a medical diagnosis dataset (tuberculosis vs. allergic sinusitis) , a derived dataset based on DDXPlus (fansi2022ddxplus). We randomly sampled a balanced subset using a fixed random seed, comprising 600 training, 100 validation, and 600 test samples, each with a 50/50 class split between tuberculosis and allergic sinusitis.
  • SPAM: an email spam detection dataset derived from deysi2023spam. We constructed a balanced subset with 600 training, 100 validation, and 600 test samples, following the same procedure as for DIAGNO.
  • JAILBREAK (shen2024donowcharacterizingevaluating): a jailbreak prompt classification dataset with 923 training, 102 validation, and 249 test samples (32 examples have been dropped due to high context tokens).
  • BANKCHURN (topre2025bankchurn): a customer–attrition prediction dataset originally provided in tabular form. Following arzaghi2025intrinsic, we serialize each record into a short textual profile describing the customer attributes. We constructed a balanced subset with 600 training, 100 validation, and 600 test samples, following the same procedure as for DIAGNO.
  • IMDB (stanford2020imdb): a sentiment analysis dataset consisting of movie reviews labeled as positive or negative. We constructed a balanced subset with 600 training, 100 validation, and 600 test samples, following the same procedure as for DIAGNO.
Appendix C Review of Traditional Decision Tree Methods: CART and C4.5
Classification and Regression trees (CART), proposed by DBLP:books/wa/BreimanFOS84, are widely used decision tree algorithms for supervised classification tasks. In this work, we focus exclusively on the classification setting.
@@ B.3 Details on Datasets
C4.5 (quinlan1993c4) follows the same recursive tree-building process as CART but differs mainly in two respects: (i) it chooses splits by maximizing the information gain ratio rather than minimizing Gini impurity, and (ii) it allows multiway splits on categorical attributes, whereas CART restricts all splits to be binary.
@@
Appendix D Additional Qualitative Results
D.1 Qualitative analysis of the questions generated by ACT